# Stellar-mass black holes in young massive and open stellar clusters – VI. Role of external galactic field

### v1

@@ Abstract
Young massive clusters (YMC) and open clusters (OC) are widely considered as potential environments for assembling merging binary stellar-remnant black holes (BBH) via dynamical interactions. However, such moderate mass systems (typically, ${}10^{3}{\rm M}_{\odot}-10^{4}{\rm M}_{\odot}$) are susceptible to being disrupted by the external tidal field of their host galaxies, potentially limiting their effectiveness as GW sources. In this study, I investigate the formation of BBH mergers in tidally dissolving star clusters. This is achieved with a newly computed grid consisting of 95 evolutionary model star clusters, where the clusters are subjected to a varying extent of tidal stripping. The cluster evolutions are computed with the direct N-body integrator ${\tt NBODY7}$ that includes, among others, treatments for post-Newtonian (PN) effects in compact-binary members, mass loss due to stellar evolution, formation of stellar remnants, and tidal stripping. It is found that even strong tidal stripping does not quench the formation of a black hole (BH) core inside a cluster or the formation of dynamical BBH mergers in the system. The overall properties of BBH mergers, e.g., the form of the distribution of merger delay time, primary mass, and mass ratio, and the redshift evolution of merger rate are not significantly altered by the extent of tidal stripping of the parent cluster population. Furthermore, even strongly tidally stripped clusters are capable of dynamically forming Gaia-BH-like detached BH–main-sequence-star binaries that escape into the galactic field. Limitations of the present study and potential future improvements are discussed.
@@ I Introduction
With mounting gravitational-wave (hereafter GW) events having been and being detected by ground-based GW interferometers [1], speculations regarding formation mechanisms of these events have intensified (see [2] for a review). The majority of the to-date-observed GW events are classified as binary stellar-remnant black hole (hereafter BBH) mergers. A diverse range of scenarios have been proposed as potential formation channels for BBH mergers. Such scenarios include dynamical interactions among stellar-remnant black holes (hereafter BH) in dense star clusters [3, 4, 5], isolated evolution of massive-stellar binaries (e.g., [6]), interacting BHs in AGN gas disks (e.g., [7, 8]), isolated evolution of hierarchical massive-stellar systems (e.g., a massive field triple-star; [9]), close flyby interactions on BBHs in the galactic field (e.g., [10]), and cluster and galactic tides on wide BBHs (e.g., [11, 12]). Despite being among the most extensively studied scenarios, there are aspects that warrant further exploration in the physics of dynamical interactions of BHs inside dense stellar systems.
@@ II.1.2 Direct, star-by-star, post-Newtonian N-body simulations
The stars in the initial models are taken to be zero age main sequence (hereafter ZAMS) stars of masses $0.08{\rm M}_{\odot}\leq m_{\ast}\leq 150.0{\rm M}_{\odot}$ that are distributed according to the Kroupa initial mass function (hereafter IMF) [69]. In all initial models, the stars have an overall (see below) primordial-binary fraction of $f_{\rm bin}=10$%. However, the initial binary fraction of the O-type stars ($m_{\ast}\geq 16.0{\rm M}_{\odot}$), which are initially paired only among themselves, is taken to be $f_{\rm Obin}(0)=100$%, which is consistent with the observed high binary fraction among O-stars in young clusters and associations [e.g. 70, 71, 72]. The O-star binaries are considered to initially follow the observed orbital-period distribution of Ref. [70] and a uniform mass-ratio distribution. The initial orbital periods of the non-O-star primordial binaries follow the period distribution of Ref. [73] and their mass-ratio distribution is also uniform. The initial eccentricity of the O-star binaries follows the eccentricity distribution of Ref. [70] and that for the rest of the binaries follows the thermal eccentricity distribution [13]. As explained in Ref. [74], such a scheme for including primordial binaries provides a reasonable compromise between the economy of computing and consistencies with observations.
@@
In the present computations, small birth spins, as per the ‘MESA’ BH-spin model of Ref. [75], are assigned to all BHs derived from single stars or from members of non-mass-transferring or non-interacting binaries. In the event of a BH-star merger (the formation of a BH Thorne–Zytkow object or BH-TZO, [76]), $f_{\rm TZ}=95$% of the merging star’s mass is assumed to be accreted onto the BH. In star-star mergers, $f_{\rm mrg}=20$% of the secondary’s mass ($\leq 10$% of the total merging stellar mass) is assumed to be lost in the merger process [77, 78]. While the extent of mass accretion onto a BH and mass loss in star-star mergers is ambiguous, the above choices favour the formation of massive BHs in moderately dense environments, especially within the PSN mass gap [79, 80]. As demonstrated in Ref. [81], that way the extent of massive BH formation in star clusters is consistent with the BBH merger rate in the PSN gap as observed by LVK [82, 83]. The above assumptions make the present N-body computations consistent with those in Ref. [81].
@@
In the present model grid, all clusters are initially tidally under-filled but the tidal filling increases (i.e., $r_{t}$ decreases) with increasing strength of the external field, i.e., with decreasing $r_{g}$ (see Eqn. 1; below). As a cluster evolves, it initially expands due to BH heating (see Sec. I), stellar-evolutionary mass loss in stellar wind, SN and planetary-nebula ejecta, and member loss due to relaxation-driven evaporation and dynamical ejections [e.g., 22, 59, 24]. As the scale length (half-mass radius, $r_{\rm h}$) of the cluster approaches $r_{t}$, the cluster starts to become tidally stripped [46], when its membership, $N$, and, hence, its relaxation time begin to decrease drastically and the latter approaches the dynamical time [13]. At the same time, the cluster deviates significantly from spherical symmetry, being bulged towards the first and the second Lagrangian points. Due to explicit, star-by-star trajectory integration without any symmetry or equilibrium assumption, ${\tt NBODY7}$ (as well other NBODYx-series codes) can reliably and uniquely handle such an extreme evolutionary condition. Most of the models in the present grid that have reached such a tidally stripping phase could be evolved until the complete dissolution ($N\lesssim 50$).
@@ III.1 General properties of cluster evolution
In general, at a given $r_{g}$ the lower is the metallicity the larger is the extent of expansion of the cluster (see Fig. 12, middle column). This is due to the fact that lower-$Z$ stars produce more massive BHs, making BH-heating more efficient, as noted in several earlier studies [e.g., 27, 24]. For a given $Z$, a cluster generally expands to a less extent when it is subjected to a stronger external field (by placing it at a smaller $r_{g}$). This is caused by enhanced tidal stripping and decay of the tidal boundary (i.e., $r_{t}$) with increasing external field.
@@ III.2 Formation of gravitational-wave mergers
Fig. 8 (upper panels) shows the present-day-observable, intrinsic distributions of the merger primary mass, ${\rm M}_{1}$, and mass ratio, $q$, as obtained from the entire model grid. Here, three cases are considered, namely, mergers from (a) only the clusters at $r_{g}>4$ kpc that remain tidally under-filled throughout their evolution and survive the tidal dissolution (Sec. III.1; Fig. 6), (b) only the clusters at $r_{g}\leq 4$ kpc that fill their tidal radii in the course of their evolution and dissolve in $<10$ Gyr due to tidal stripping (the tidally stripped case), and (c) clusters with $r_{g}$ chosen randomly within $1.0{\rm~{}kpc}\leq r_{g}\leq 8.5{\rm~{}kpc}$ (the random tidal filling case). Over the respective $r_{g}$-ranges, the cluster $r_{g}$s and $r_{\rm h}(0)$s are chosen randomly. The present-day-observable distributions are obtained by applying the Monte Carlo approach as detailed in Ref. [92], which description, for brevity, is not repeated here. As in the above reference, the cosmic SFR evolution as of Ref. [93] (up to redshift $z=10$), the redshift-metallicity dependence as of Ref. [94] (their ‘moderate-$Z$’ dependence), a GW event visibility horizon of $z_{\rm max}=1$, and the $\Lambda$CDM cosmological framework [95, 96, 97] with the cosmological constants from Planck ($H_{0}=67.4{\rm~{}km~{}s}^{-1}{\rm~{}Mpc}^{-1}$, $\Omega_{\rm~{}m}=0.315$, flat Universe for which $t_{\rm Hubble}=13.79{\rm~{}Gyr}$ [98]) are adopted. A total of $10^{6}$ clusters are drawn from the model grid for the Monte Carlo sampling [92], in each case. Since, in this study, only one cluster mass is considered, only the normalized distributions (probability density function or PDF) are plotted in Fig. 8 (upper panels) and not the total (differential) merger rate distributions. That way, Fig. 8 (upper panels) shows only the forms of the distributions in the different cases and the differences between them. As seen, the ${\rm M}_{1}$ and $q$ distributions are overall similar in form for the three cases. A few mergers occur within the PSN gap (i.e., with ${\rm M}_{1}>40{\rm M}_{\odot}$), mainly due to the formation of over-massive BHs via star-star mergers [99, 32, 100, 81].
@@ III.3 Formation of black hole-main sequence binaries
The current models also produce detached (i.e., non-mass-transferring) BH-star binaries that get ejected from the cluster and escape to the galactic field. The formation mechanisms and the overall properties of these field BH-star binaries are similar to those found in earlier star cluster computations by the author. The discussion is, therefore, not repeated here and the reader is encouraged to consult Ref. [101] for all the details. Fig. 10 shows the ejected black hole-main sequence (hereafter BH-MS) binaries from only the present model grid, distinguishing between dynamically paired and primordially paired systems. Consistently with the previous work (see also Refs. [102, 103]), the parameters of the dynamically paired field BH-MS binaries from the clusters encompass those of the observed, detached BH-MS binary candidates Gaia-BH1 [104] and Gaia-BH3 [105] in the MW field. However, the parameters of the primordially paired ones disagree rather strongly with these observed binaries’ parameters. For Gaia-BH1, the disagreement is mainly due to its low component-star mass that the primordially paired BH-MS binaries hardly possess (see Fig. 10). For Gaia-BH3, the disagreement with the primordially paired population is due to the binary’s low component-star mass, high BH mass, and long orbital period.
The total formation efficiency of field BH-MS binaries containing a low mass ($<3{\rm M}_{\odot}$) MS companion from the present models is $8.4\times 10^{-6}{\rm M}_{\odot}^{-1}$. Restricting to only solar-mass-like ($0.7{\rm M}_{\odot}-1.1{\rm M}_{\odot}$) companions, the formation efficiency is $2.1\times 10^{-6}{\rm M}_{\odot}^{-1}$. All of the escaped BH-MS binaries containing a low mass MS companion are dynamically paired (c.f. Fig. 10). As discussed in Ref. [101], the rather low formation efficiency of escaped BH-star binaries is due to the fact that the dynamically interacting BH-star systems inside the clusters favourably undergo exchange interactions to become a BBH.
@@ IV Summary and discussions
Notably, some recent authors have considered star cluster models with extremely compact, $\approx 0.1-0.3$ pc, initial sizes (e.g., [107, 106, 37]). By evolving from such a compact initial size, a cluster would potentially be much less affected by the external field. However, such initial scales represent those of gas-embedded clusters [40]. It can be expected that the gas eventually gets expelled due to radiation pressure and mechanical (e.g., wind) feedback from the cluster’s massive stars [108]. With a typical $\approx 30$% star-to-gas fraction [109], or with even smaller star formation efficiency [110, 111], the gas expulsion would expand the cluster to a pc-scale size within a few Myr age, if not dissolve the cluster completely [112, 109, 39]. However, there is ambiguity regarding whether gas can always be successfully expelled from a newborn embedded cluster, or whether it is possible to convert all of the gas into stars [113, 114, 39]. Observationally, YMCs and OCs are typically gas free and are of parsecs length-scale [38], consistent with the initial conditions adopted here.
@@
It is nevertheless necessary to study the effect of galactic tidal field with additional cluster masses (both lower and higher masses). This will help to convolve the effect with a cluster mass function. It would also be worthwhile to consider cluster evolution under the influence of the potential of an irregular and asymmetrical field mass distribution, such as those obtained from high-resolution structure formation simulations (e.g., [115, 116]). Such studies will be taken up in the near future.

### v2

@@ Abstract
Young massive clusters (YMC) and open clusters (OC) are widely considered as potential environments for assembling merging binary stellar-remnant black holes (BBH) via dynamical interactions. However, such moderate mass systems (typically, $10^{3}{\rm M}_{\odot}-10^{4}{\rm M}_{\odot}$) are susceptible to being disrupted by the external tidal field of their host galaxies, potentially limiting their effectiveness as gravitational-wave (GW) sources. In this study, I investigate the formation of BBH mergers in tidally dissolving star clusters. This is achieved with a newly computed grid consisting of 95 evolutionary model star clusters, where the clusters are subjected to a varying extent of tidal stripping. The cluster evolutions are computed with the direct N-body integrator $\tt NBODY7$} that includes, among others, treatments for post-Newtonian (PN) effects in compact-binary members, mass loss due to stellar evolution, formation of stellar remnants, and tidal stripping. It is found that even strong tidal stripping does not quench the formation of a black hole (BH) core inside a cluster or the formation of dynamical BBH mergers in the system. The overall properties of BBH mergers, e.g., the form of the distribution of merger delay time, primary mass, and mass ratio, and the redshift evolution of merger rate are not significantly altered by the extent of tidal stripping of the parent cluster population. Furthermore, even strongly tidally stripped clusters are capable of dynamically forming Gaia-BH-like detached BH–main-sequence-star binaries that escape into the galactic field. Limitations of the present study and potential future improvements are discussed.
@@ I Introduction
With mounting gravitational-wave (hereafter GW) events having been and being detected by ground-based GW interferometers [1], speculations regarding formation mechanisms of these events have intensified (see Refs. [2, 3] for comprehensive reviews). The majority of the to-date-observed GW events are classified as binary stellar-remnant black hole (hereafter BBH) mergers. A diverse range of scenarios have been proposed as potential formation channels for BBH mergers. Such scenarios include dynamical interactions among stellar-remnant black holes (hereafter BH) in dense star clusters [4, 5, 6], isolated evolution of massive-stellar binaries (e.g., [7]), interacting BHs in AGN gas disks (e.g., [8, 9]), isolated evolution of hierarchical massive-stellar systems (e.g., a massive field triple-star; [10]), close flyby interactions on BBHs in the galactic field (e.g., [11]), and cluster and galactic tides on wide BBHs (e.g., [12, 13]). Despite being among the most extensively studied scenarios, there are aspects that warrant further exploration in the physics of dynamical interactions of BHs inside dense stellar systems.
@@ II.1.2 Direct, star-by-star, post-Newtonian N-body simulations
The stars in the initial models are taken to be zero age main sequence (hereafter ZAMS) stars of masses $0.08{\rm M}_{\odot}\leq m_{\ast}\leq 150.0{\rm M}_{\odot}$ that are distributed according to the Kroupa initial mass function (hereafter IMF) [70]. In all initial models, the stars have an overall (see below) primordial-binary fraction of $f_{\rm bin}=10$%. However, the initial binary fraction of the O-type stars ($m_{\ast}\geq 16.0{\rm M}_{\odot}$), which are initially paired only among themselves, is taken to be $f_{\rm Obin}(0)=100$%, which is consistent with the observed high binary fraction among O-stars in young clusters and associations [e.g. 71, 72, 73]. The O-star binaries are considered to initially follow the observed orbital-period distribution of Ref. [71] and a uniform mass-ratio distribution. The initial orbital periods of the non-O-star primordial binaries follow the period distribution of Ref. [74] and their mass-ratio distribution is also uniform. The initial eccentricity of the O-star binaries follows the eccentricity distribution of Ref. [71] and that for the rest of the binaries follows the thermal eccentricity distribution [14]. As explained in Ref. [75], such a scheme for including primordial binaries provides a reasonable compromise between the economy of computing and consistencies with observations.
@@
In the present computations, small birth spins, as per the ‘MESA’ BH-spin model of Ref. [76], are assigned to all BHs derived from single stars or from members of non-mass-transferring or non-interacting binaries. In the event of a BH-star merger (the formation of a BH Thorne–Zytkow object or BH-TZO [77]), $f_{\rm TZ}=95$% of the merging star’s mass is assumed to be accreted onto the BH. In star-star mergers, $f_{\rm mrg}=20$% of the secondary’s mass ($\leq 10$% of the total merging stellar mass) is assumed to be lost in the merger process [78, 79]. While the extent of mass accretion onto a BH and mass loss in star-star mergers is ambiguous, the above choices favour the formation of massive BHs in moderately dense environments, especially within the PSN mass gap [80, 81]. As demonstrated in Ref. [82], that way the extent of massive BH formation in star clusters is consistent with the BBH merger rate in the PSN gap as observed by LVK [83, 84]. The above assumptions make the present N-body computations consistent with those in Ref. [82].
@@
In the present model grid, all clusters are initially tidally under-filled but the tidal filling increases (i.e., $r_{t}$ decreases) with increasing strength of the external field, i.e., with decreasing $r_{g}$ (see Eqn. 1; below). As a cluster evolves, it initially expands due to BH heating (see Sec. I), stellar-evolutionary mass loss in stellar wind, SN and planetary-nebula ejecta, and member loss due to relaxation-driven evaporation and dynamical ejections [e.g., 23, 60, 25]. As the scale length (half-mass radius, $r_{\rm h}$) of the cluster approaches $r_{t}$, the cluster starts to become tidally stripped [47], when its membership, $N$, and, hence, its relaxation time begin to decrease drastically and the latter approaches the dynamical time [14]. At the same time, the cluster deviates significantly from spherical symmetry, being bulged towards the first and the second Lagrangian points. Due to explicit, star-by-star trajectory integration without any symmetry or equilibrium assumption, $\tt NBODY7$} (as well other NBODYx-series codes) can reliably and uniquely handle such an extreme evolutionary condition. Most of the models in the present grid that have reached such a tidally stripping phase could be evolved until the complete dissolution ($N\lesssim 50$).
@@ III.1 General properties of cluster evolution
In general, at a given $r_{g}$ the lower is the metallicity the larger is the extent of expansion of the cluster (see Fig. 12, middle column). This is due to the fact that lower-$Z$ stars produce more massive BHs, making BH-heating more efficient, as noted in several earlier studies [e.g., 28, 25]. For a given $Z$, a cluster generally expands to a less extent when it is subjected to a stronger external field (by placing it at a smaller $r_{g}$). This is caused by enhanced tidal stripping and decay of the tidal boundary (i.e., $r_{t}$) with increasing external field.
@@ III.2 Formation of gravitational-wave mergers
[…]
@@ III.3 Formation of black hole-main sequence binaries
The current models also produce detached (i.e., non-mass-transferring) BH-star binaries that get ejected from the cluster and escape to the galactic field. The formation mechanisms and the overall properties of these field BH-star binaries are similar to those found in earlier star cluster computations by the author. The discussion is, therefore, not repeated here and the reader is encouraged to consult Ref. [102] for all the details. Fig. 10 shows the ejected black hole-main sequence (hereafter BH-MS) binaries from only the present model grid, distinguishing between dynamically paired and primordially paired systems. Consistently with the previous work (see also Refs. [103, 104]), the parameters of the dynamically paired field BH-MS binaries from the clusters encompass those of the observed, detached BH-MS binary candidates Gaia-BH1 [105, 106] and Gaia-BH3 [107] in the MW field. However, the parameters of the primordially paired ones disagree rather strongly with these observed binaries’ parameters. For Gaia-BH1, the disagreement is mainly due to its low component-star mass that the primordially paired BH-MS binaries hardly possess (see Fig. 10). For Gaia-BH3, the disagreement with the primordially paired population is due to the binary’s low component-star mass, high BH mass, and long orbital period.
The total formation efficiency of field BH-MS binaries containing a low mass ($3{\rm M}_{\odot}$) MS companion from the present models is $8.4\times 10^{-6}{\rm M}_{\odot}^{-1}$. Restricting to only solar-mass-like ($0.7{\rm M}_{\odot}-1.1{\rm M}_{\odot}$) companions, the formation efficiency is $2.1\times 10^{-6}{\rm M}_{\odot}^{-1}$. All of the escaped BH-MS binaries containing a low mass MS companion are dynamically paired (c.f. Fig. 10). As discussed in Ref. [102], the rather low formation efficiency of escaped BH-star binaries is due to the fact that the dynamically interacting BH-star systems inside the clusters favourably undergo exchange interactions to become a BBH.
@@ IV Summary and discussions
Another potential drawback is the representation of the model grid with only one evolutionary model cluster per grid point. At present, this limitation stems primarily from the high computational cost of direct N-body simulation of models initiating with $M_{\rm cl}(0)\approx 3\times 10^{4}{\rm M}_{\odot}$ (membership $N(0)\approx 5\times 10^{4}$) and containing primordial binaries (see Sec. II.1.2). As each model cluster typically produces a few GR mergers, the mergers’ delay times and rate distributions are prone to statistical uncertainties that are not represented in the current results. Qualitative agreements in trends are nevertheless apparent between the different tidal-filling cases and metallicities considered (Figs. 8 and 9).
Notably, a model cluster grid that includes variation in $r_{g}$ in a MW-like external potential has also been presented in Ref. [25]. This work comprises Monte Carlo N-body evolutionary simulations [15] of a grid of GC-like model clusters (the ‘CMC Cluster Catalogue’) that are $\sim 10$ times more massive than the present models. Expectedly, although these models do exhibit a stronger disruption with decreasing $r_{g}$ (see their Table A1), they are generally more resilient to tidal disruption. Accordingly, these models do not show any particular trend in the BBH-mergers (and compact binary formation) from them (see their Table A5). It is, however, noteworthy that CMC model computations (Monte Carlo N-body simulations in general) need to be halted well before the two-body relaxation timescale of a cluster becomes comparable to the cluster’s dynamical timescale (see Ref. [109] and references therein) as the cluster continues to deplete. That way, the signatures of enhanced tidal stripping may not be fully captured in the CMC models.
The BBH-merger formation efficiencies of the CMC cluster models are typically $\sim 10^{-4}{\rm M}_{\odot}^{-1}$, i.e., $\sim 10$ times higher than those of the current models. This difference is likely due to the much higher, GC-like masses of the CMC models and also due to their assumption of the ‘rapid’ remnant mass model, as opposed to the delayed model in the present computations. The rapid model causes a higher BH retention in clusters than its delayed counterpart [33]. Interestingly, more recent direct N-body computations of star clusters of masses comparable to the CMC models (the ‘DRAGON-II’ simulations [32]) yield merger efficiencies $\sim 10^{-5}{\rm M}_{\odot}^{-1}$, i.e., similar to those of the present models.
@@
Recently, several authors have considered star cluster models with extremely compact, $\approx 0.1-0.3$ pc, initial sizes (e.g., Refs. [110, 108, 38]). By evolving from such a compact initial size, a cluster would potentially be much less affected by the external field. However, such initial scales represent those of gas-embedded clusters [41]. It can be expected that the gas eventually gets expelled due to radiation pressure and mechanical (e.g., wind) feedback from the cluster’s massive stars [111]. With a typical $\approx 30$% star-to-gas fraction [112], or with even smaller star formation efficiency [113, 114], the gas expulsion would expand the cluster to a pc-scale size within a few Myr age, if not dissolve the cluster completely [115, 112, 40]. However, there is ambiguity regarding whether gas can always be successfully expelled from a newborn embedded cluster, or whether it is possible to convert all of the gas into stars [116, 117, 40]. Observationally, YMCs and OCs are typically gas free and are of parsecs length-scale [39], consistent with the initial conditions adopted here.
Furthermore, star clusters are observed to be born inside dense molecular clouds, where the star formation is heavily sub-structured and/or distributed along gas filaments [118, 41]. It is widely conjectured that the present-day-observed, near monolithic YMCs are hierarchically assembled from such an irregular, gas-embedded configuration by undergoing a violent relaxation phase. A YMC is formed if the stellar system survives the violent relaxation, gas dispersal, and potential stripping by external tidal field [116, 40]. The monolithic, virial, gas-free initial cluster configuration, as adopted here, therefore implicitly preselects those systems that survive the cluster assembly phase. This corresponds to an effective cluster formation efficiency (hereafter CFE) for YMCs, whose estimate is rather ambiguous in the literature [119, 120]. Merger rate estimates are dependent on the CFE, but since relative rate distributions are mainly considered in this work (see Fig. 8 and the associated description), the CFE value is irrelevant for the present results. Also, $\eta_{\rm mrg}$ is, by definition, independent of CFE. The present results, however, implicitly assume that the CFE remains constant with redshift and $r_{g}$, as commonly adopted in the literature due to the lack of estimates of such dependencies.
The early hierarchical assembly phase would cause an enhanced mass segregation at the young age of the cluster [121, 122]. However, since the mass segregation timescale of the BH-progenitor stars in such clusters is typically $\lesssim 1$ Myr [14, 18], an initially mass-segregated state of the models (as opposed to the presently assumed initially unsegregated state) would unlikely cause qualitative differences in the clusters’ evolution and their BBH-merger outcomes. The assembly phase can also potentially cause a runaway growth of an intermediate-mass black hole, but this process is probable in a significantly more massive stellar system [123, 124]. In summary, it can be expected that the complex star formation morphology and assembly phase of YMCs are unlikely to qualitatively alter the current inferences.
@@
It is nevertheless necessary to study the effect of galactic tidal field with additional cluster masses (both lower and higher masses). This will help to convolve the effect with a cluster mass function. It would also be worthwhile to consider cluster evolution under the influence of the potential of an irregular and asymmetrical field mass distribution, such as those obtained from high-resolution structure formation simulations (e.g., Refs. [125, 126]). Such studies will be taken up in the near future.